OpenStudy (anonymous):

If f(x)=5x-3, then what is f(x-4)?

OpenStudy (anonymous):

Simply plug x-4 into your equation and simplify

OpenStudy (anonymous):

f(x)=5(x)-3 f(x-4)=5(x-4)-3

OpenStudy (anonymous):

Thanks